OpenStudy (anonymous):

8y - 4 + 3(y + 7) = 6y - 3(y - 3)

OpenStudy (anonymous):

8y-4+3y+21=6y-3y+9 11y+17=3y+9 8y=-8 y=-1

OpenStudy (anonymous):

i see thnkz alot

OpenStudy (anonymous):

np